OpenStudy (anonymous):

Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2

OpenStudy (anonymous):

slope is \(\frac{9-5}{x-10}\) and you want that to be equal 2, so set \[\frac{4}{x-10}=2\] and solve for \(x\)

OpenStudy (anonymous):

you get \[\frac{4}{x-10}=2\] \[4=2(x-10)\] \[2=x-10\] \[x=12\]